Assume the risk free rate is 2.0%. The SP500 is considered the market. Today (T-0), you invest $400 in Stock A and $600 in Stock B to create Portfolio A,B. Assume there are not taxes or dividends. The one year performance of the stocks and the market is summarized in the table below. Use this information to help answer questions 16-20 Investment Total Return 12.0% 18.0% 10.0% Total Risk 14.0% 15.0% 12.0% Beta 1.00 1.40 0.60 Market Stock A 400 Stock B 600 16. Portfolio A,Bs beta is closest to: a. 0.80 b. 0.92 c. 1.00 d. 1.08 e. 1.20 17. If you doubled the size of your investment today (T-O), your total dollar return would be closest to: a. $66 b. $92 c. $184 d. $264 e. $280 18. Assuming the correlation of stock A & Bis zero (which your book assumes for all securities), Portfolio A,Bs total risk is closest to: 94% 10.2% b. d. 12.8% 13.2% e. 19. Compare the SP500, Stock A, Stock B, and Portfolio A,B. Which of the following statements is most likely TRUE: a. Stocks B Share Ratio of 0.67 is better than the Sharpe Ratio of the Stock A, SP500, and Portfolio A,B b. Stocks A Share Ratio of 1.07 is better than the Sharpe Ratio of the Stock B, SP500, and Portfolio A,B c. Stocks A Share Ratio of 1.20 is better than the Sharpe Ratio of the Stock B, SP 500, and Portfolio A,B d. Stocks As Share Ratio of 1.40 is better than the Sharpe Ratio of the Stock B, SP500, and Portfolio A,B e. Portfolio A,Bs Sharpe Ratio of 1.20 is better than the Sharpe Ratio of the Stock A, Stock B, and the SP500 20. In lecture we discussed how assuming the correlation between any two is zero is dangerous when assessing the total risk of a portfolio between two assets to be 0.40 instead of 0.00. Portfolio A,Bs total risk is now closest to: roble assume you calculate the correlation 9.4% 10.2% о. b. d. 12.8% 13.2% e.
Answer #1

16) Portfolio beta=(wt of stock A*beta of A)+(wt of stock B*beta B)
wt of stock A= investment in A/total investment
=400/(400+600)=40%
wt of stock B= investment in B/total investment
=600/(400+600)=60%
Portfolio Beta=(0.4*1.4)+(0.6*0.6)=0.92
It is option B

17) since we doubling our investment amiunt invested in each and return is below
stock A=400*2=800 and return=800*18%=144
stock B=600*2=1200 and return=1200*10%=120
total return=144+120=264
it is option D

18)std of portfolio=((wt of A* std of A)^2+(wt of B* std of B)^2+(2* wt of A*wt of B* std A*std B* correlation(A,B))^0.5
=[((0.4*15%)^0.5)+((0.6*12%)^0.5)]^0.5
=9.4%
it is option A

19)Sharpe ratio=(return stock-risk free)/std of stock
Portfolio return=(0.4*18%)+(0.6*10%)=13.2%
Sharpe ratio A=(18%-2%)/15%=1.07
Sharpe ratio B=(10%-2%)/12%=0.67
sharpe ratio portfolio A,B=(13.2%-2%)/9.4%=1.19
SP 500 =(12%-2%)/14%=0.71
It is option E

20)std of portfolio=((wt of A* std of A)^2+(wt of B* std of B)^2+(2* wt of A*wt of B* std A*std B* correlation(A,B))^0.5
=[((0.4*15%)^0.5)+((0.6*12%)^0.5)]^0.5+(2*0.4*0.6*15%*12%*0.4)'^0.5
=11.1%
It is option C